No textbook, came from a worksheet
A. Slope = -2, through (2,3)
B. m = -2/3, throught (-3,2)
C. through (6,3) and (5,4)

Answer by stanbon(75887)

For each problem you need an equation of the form y=mx+b
------------------------------------------
A. Slope = -2, through (2,3)
You are given x=2,y=3, m=-2. You need to find "b"
3=(-2)2+b
b=3+4=7
EQUATION: y=-2x+7
B. m = -2/3, throught (-3,2)
Same proceudure here because you are given x,y, and m
C. through (6,3) and (5,4)
Here you have to find m:
m=[4-3]/[5-6]=-1
Now you have x=6,y=3,m=-1
Proceed as I did for problem "A".
Cheers,
Stan H.